OpenStudy (anonymous):

i give metals Algebra 2

OpenStudy (anonymous):

Compare the functions shown below: f(x) = 4 sin (2x –π) - 1 g(x) x y -1 6 0 1 1 -2 2 -3 3 -2 4 1 5 6 h(x) = (x –2)2 + 4 Which function has the smallest minimum y-value?

OpenStudy (anonymous):

@Directrix @satellite73 @DNTU4GETBOUTME

OpenStudy (anonymous):

f(x)=ax2+ b x+ cf(x)=3x2+12x+16

OpenStudy (anonymous):

looks like the minimum value of g is -3

OpenStudy (anonymous):

the minimum values of \((x-2)^2+4\) must be \(4\) right ?

OpenStudy (anonymous):

the minimum value of sine is -1, so the minimum value of 4 sine is -4, making the minimum value of 4 sin -1 equal -5